OpenStudy (anonymous):

Simplify: 9√7 - √28 a. 5√7 b. 7√7 c. 11√7 d. 13√7

hero (hero):

11 root 7

OpenStudy (anonymous):

eh? √28 = √4 * √7 = 2√7 9√7 - 2√7 = 7√7 b